# Investigating and Mitigating Stereotype-aware Unfairness in LLM-based Recommendations

### v1

@@ 4.1.2. Evaluation Metrics
MoS (Ours) two-sided 55.71(-5.0%) 50.04(-16.6%) 57.24(-3.6%) 79.15(+1.1%) 79.39(-9.3%) 72.98(+22.9%)
  • ::*:∗ : The term $\mathcal{L}_{\mathrm{fair}}$ in $\mathcal{L}$ according to Eq. (14) is replaced by corresponding fairness metrics proposed in each baseline of fairness-oriented methods (Jiang et al., 2024; Hua et al., 2023a; Zhang et al., 2023).
Table 5. Results of LLM-RS paradigms. We report the average results over three independent runs with random seeds. The reported results are multiplied by 100, where boldface and underline indicate the best and second best score, respectively.
@@
MovieLens TALLRec 0.3954 $\mathcal{L}_{\mathrm{rec}}$ 77.05 80.94 74.99 70.27 68.32 86.33
${}^{\text{-{}-}}$ 0.0138 $\mathcal{L}_{\mathrm{rec}}$ 80.00 81.59 78.57 65.57 67.62 80.71
MoS (Ours) 0.0331 $\mathcal{L}$ 55.49 53.71 50.91 69.64 77.83 50.80
- prompt-tuning 0.0138 $\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{fair}}$ 65.53 66.25 61.43 56.76 67.65 44.05
- p-tuning 0.1097 $\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{fair}}$ 67.60 72.07 60.49 67.23 72.90 55.79
- LoRA 0.3954 $\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{fair}}$ 58.46 63.18 60.49 68.80 72.06 75.88
BookCrossing TALLRec 0.3954 $\mathcal{L}_{\mathrm{rec}}$ 59.48 59.65 58.61 79.43 70.85 90.12
${}^{\text{-{}-}}$ 0.0138 $\mathcal{L}_{\mathrm{rec}}$ 58.67 61.25 64.57 79.57 74.19 83.36
MoS (Ours) 0.0331 $\mathcal{L}$ 55.71 50.04 57.24 79.15 79.39 72.98
- prompt-tuning 0.0138 $\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{fair}}$ 54.07 60.30 60.56 78.34 88.18 43.00
- p-tuning 0.1097 $\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{fair}}$ 58.35 63.65 64.29 79.01 77.49 74.30
- LoRA 0.3954 $\mathcal{L}_{\mathrm{rec}}+\mathcal{L}_{\mathrm{fair}}$ 56.25 58.67 59.48 79.56 85.51 60.30
[…]
Performance. Following the implementation of TALLRec, we adopt the area under the receiver operating characteristic (AUC) for performance evaluation. In addition, we compare the Precision and Recall to probe over-recommendations (i.e., false positive predictions) between different stereotype groups.
4.1.3. Baselines
@@
Fairness-oriented Methods. IFairLRS (Jiang et al., 2024) proposes a reweighting strategy to mitigate biases stemming from unbalanced groups, where sample weights are applied to the loss of instruction-tuning samples of LLM-RS. UP5 (Hua et al., 2023a) introduces a counterfactually fair prompting method, which masks sensitive user information by prompt tuning with a discrimination loss. FaiRLLM (Zhang et al., 2023) designs fairness metrics tailored to LLM-RS by mitigating the divergence of similarity metrics of recommendations against sensitive attributes. LLM-RS Paradigms. TALLRec (Bao et al., 2023) proposes a lightweight and effective paradigm to adapt LLMs to recommendation tasks with PEFT, which serves as our baseline of LLM-RS. To assess both the recommendation performance and the fairness of mitigating stereotypes in LLM-RS, we further compare our proposed MoS framework for training LLM-RS with conventional PEFT paradigms, including prompt tuning (Lester et al., 2021), p-tuning (Liu et al., 2022), and LoRA (Hu et al., [n. d.]).
4.1.4. Implementation Details
To implement LLM-RS in a generative paradigm, the recommendation task is formulated into prompt. In particular, the personalized preference of users is indicated based on median ratings, which are 3 and 5 in MovieLens and BookCrossing datasets, respectively. An example prompt is provided below:
@@ 4.1.4. Implementation Details
Our proposed methods are implemented based on HuggingFace and PyTorch. For a fair comparison across all baselines, we employ a widely-used lightweight LLM, i.e., T5 (Raffel et al., 2020), with encoder-decoder structures as the backbone model of LLM-RS. As for the proposed MoS framework, we employ linear stereotype-wise routing models (i.e., router and reweighting layer) and $4$ prompt-based experts, each with a length of $5$ tokens for eliciting different stereotypes in personalized recommendations. In particular, we optimize the aforementioned proposed models with Adafator (Shazeer and Stern, 2018), where the learning rates for prompt-based baselines (i.e., p-tuning, prompt tuning, and MoS) and adapter-based baselines (i.e., TALLRec and LoRA) are set to be $0.5$ and $0.005$, respectively.
4.2. Performance Comparison
4.2.1. Comparison of Fairness-oriented Methods
As shown in Table 4, we compare the fairness and performance between our proposed MoS and existing fairness-oriented methods of LLM-RS. Overall, MoS significantly outperforms the current single-sided fairness baselines (i.e., user-side and item-side) to mitigate stereotypes in LLM-RS. In the meanwhile, MoS achieves slightly better or comparable recommendation performance compared to each baseline.
@@ 4.2.1. Comparison of Fairness-oriented Methods
4.2.2. MoS vs. Conventional LLM-RS Paradigms
To demonstrate the effectiveness of MoS to learn unbiased representations utilizing multiple stereotype-relevant experts, we design baselines by replacing the MoS framework (i.e., multiple experts with stereotype-wise routing) with conventional PEFT methods to train LLM-RS under various setups of learning objectives, as shown in Table 5.
Comparing the fairness performance of training LLM-RS with the proposed MoS framework to conventional PEFT paradigms (i.e., prompt tuning, p-tuning, and LoRA), it can be noticed that MoS consistently outperforms all baselines by 2%-15% in fairness and achieves comparable performance between -0.9% and 18% compared to baselines even with higher trainable parameters of 331%-1194%. Notably, significant trade-offs between fairness and performance can be noticed in some conventional LLM-RS paradigms, while MoS maintains the best or second best scores of both fairness and performance in most situations.
4.3. Ablation Study
4.3.1. MoS Components
To assess the influence of each key component, we conducted ablation experiments on the effectiveness of MoS with separately eliminated components, as shown in Table 6. We compare the stereotype-wise routing component between top-1, top-2, and stochastic routing strategies, as illustrated in Eq. (7). Notably, the top-1 setting outperforms top-2 and stochastic settings in terms of both fairness and performance. We speculate that the top-2 routing strategy potentially encodes different stereotype information to the same experts. In addition, we compare the effectiveness of MoS without the reweighting layer and corresponding learning objectives $\mathcal{L}_{div}$. Despite comparable results in performance, the proposed components significantly improve the fairness by 3%-16%.
4.3.2. User-side vs. Item-side Stereotypes
Since a stereotype group contains both users and items in LLM-RS, we further delve into the effectiveness of MoS by separately comparing user-side stereotype and item-side stereotypes, as shown in Figure 4. By comparing between red/blue and yellow/green bars, it can be observed that MoS can achieve consistent fairness and performance between users and items of different stereotype groups. However, the fairness and performance of user stereotype group $u\in G_{1}$ (i.e., red and yellow bars) significantly exceed that of $u\in G_{2}$ (i.e., blue and green bars). One likely reason is that the amount of training data is dominated by $u\in G_{1}$ due to the unequal distribution of user-item interactions in recommendation datasets, as shown in Table 3.
4.3.3. Z-score for Stereotype Measurement
To validate the threshold of stereotype measurement based on Z-scores, we compare the recommendation performance of target items below and above the designed threshold. As revealed by our preliminary experiments in Table 1, items below the threshold are weakly biased to any stereotype group, indicating a lower degree of stereotype (i.e., $d_{\hat{v}\in G}$) that potentially downgrades LLM-RS performance. As shown in Figure 5, the recommendation performance of items below the threshold notably exceeds that of items above the threshold, implying that Z-scores can identify an effective threshold of stereotype measurement.
Table 6. Results of ablation studies on MoS components. We report the average fairness ($\bm{SF}$) and performance (AUC) over three independent runs with random seeds.
@@ 4.3.3. Z-score for Stereotype Measurement
Figure 4. Results of ablation studies on user-side and item-side stereotypes. The detailed statistics of users $u$, items $v$, and groups $G_{1},G_{2}$ can be found in Table 3.
5. RELATED WORK
In this section, we first review the existence and cause of intrinsic stereotypes in LLMs, then move to the existing research on the fairness issues in LLM-based recommendations.
5.1. Stereotypes of LLMs
Stereotypes refer to social bias and discrimination that associate diverse combinations of characteristics with specific groups (Kotek et al., 2023). Existing studies have revealed that LLMs trained on large-scale uncurated data, particularly biased text corpus, inherit various stereotypes against specific social groups, which can be categorized by age, gender, and religion, etc. (Navigli et al., 2023; Gallegos et al., 2024). More specifically, such LLM-encoded stereotypes are encapsulated in word embeddings, exhibiting stereotypical behavior due to the ubiquitous existence of text (e.g., prompt) in LLM-based applications (Guo et al., 2022). Compared to conventional recommendation models, LLMs introduce unique stereotypes for inferring user preferences, leading to substantial discrimination in personalized recommendations. In other words, discriminative predictions can be elicited by prompting LLMs with text-based representations of users and items. For example, a ”female” stereotype can be probed by LLMs given an interacted item ”nurse” in job recommendations (Xu et al., 2023).
@@ 5.1. Stereotypes of LLMs
Figure 5. Results of ablation studies on the threshold of stereotype measurement based on Z-scores.
5.2. Fairness in LLM-based Recommendations
Most existing studies on LLM-RS fairness focus on either user-side fairness to achieve consistent recommendation performance across user groups (Hua et al., 2023a; Zhang et al., 2023; Deldjoo and Di Noia, 2024) or item-side fairness by providing fair exposure opportunities across item groups (Jiang et al., 2024; Bao et al., 2024). However, stereotypes are embedded ubiquitously in word embeddings of LLM-RS, which simultaneously affect user-side and item-side fairness. In addition, current fairness methods might fall short in addressing stereotypes in prompts, such as item titles and descriptions, by merely leveraging discrete IDs or ID embeddings of users and items for recommendations (Hua et al., 2023a; Wang et al., 2024). In other words, such fairness methods implicitly bypass stereotypes encoded in LLM-RS. Notably, recent studies have indicated that textual knowledge is critical for harnessing the linguistic capabilities of LLM-RS to effectively comprehend user preferences in recommendations (Liao et al., 2023; Zhu et al., 2024). To sum up, effective methods to tackle the fairness against intrinsic stereotypes in LLM-RS, specifically against biased textual knowledge, remain underexplored.
6. CONCLUSION
@@
This study investigates the unique characteristics of stereotypes in LLM-RS that simultaneously exist in the word embeddings of users and items. In this paper, we propose a new variant of fairness between stereotype groups containing both users and items (i.e., two-sided groups) in LLM-RS, rather than separately considering user groups or item groups. To mitigate unfairness due to stereotypes in LLM-RS, a novel framework called MoS is proposed along with effective learning objectives. In particular, we develop multiple stereotype-relevant experts to capture different stereotypes in textual user and item information, where an insightful stereotype-wise routing strategy is designed to learn unbiased representations against different stereotypes over multiple stereotype-relevant experts. Through comprehensive experiments on recommendation datasets under various fairness settings, we demonstrated the effectiveness of the proposed methods in addressing stereotype-aware fairness of LLM-RS. As for future work, further investigation might shed light on the individual-level fairness against diverse combinations of stereotypes in LLM-RS.

### v2

@@ VII-A2 Evaluation Metrics
MoS two-sided 55.71(5.0%) 50.04(16.6%) 57.24(3.6%) 79.15(1.1%) 79.39(9.3%) 72.98(22.9%)
  • ::*:∗ : The learning objectives $\mathcal{L}_{\mathrm{fair}}$ in Eq. (15) are replaced by the corresponding fairness metrics in each baseline of fairness-oriented methods.
TABLE V: Results of LLM-RS paradigms. We report the average results over three independent runs with random seeds. The reported results are multiplied by 100, where boldface and underline indicate the best and second best score, respectively.
@@
MovieLens Full-model tuning - $\mathcal{L}_{\mathrm{rec}}$ 75.22 79.17 72.09 67.16 75.83 69.62
MoS 0.0331 72.47 80.26 70.45 65.86 69.07 64.69
Full-model tuning - $\mathcal{L}_{\mathrm{total}}*$ 62.39 62.83 61.12 68.91 73.12 62.20
MoS (multi-expert) 0.0331 $\mathcal{L}_{\mathrm{total}}$ 55.49 53.71 50.91 69.64 77.83 50.80
- LoRA (adapter) 0.3954 $\mathcal{L}_{\mathrm{total}}*$ 63.25 64.53 61.27 68.25 72.55 70.29
- p-tuning (encoder) 0.1097 $\mathcal{L}_{\mathrm{total}}*$ 67.60 72.07 60.49 67.23 72.90 55.79
- prompt-tuning (embedding) 0.0138 $\mathcal{L}_{\mathrm{total}}*$ 65.53 66.25 61.43 56.76 67.65 44.05
BookCrossing Full-model tuning - $\mathcal{L}_{\mathrm{rec}}$ 59.10 62.39 59.28 77.59 80.62 64.14
MoS 0.0331 58.25 60.02 59.45 77.22 76.45 68.18
Full-model tuning - $\mathcal{L}_{\mathrm{total}}*$ 56.14 58.67 60.96 78.41 84.25 68.05
MoS (multi-expert) 0.0331 $\mathcal{L}_{\mathrm{total}}$ 55.71 50.04 57.24 79.15 79.39 72.98
- LoRA (adapter) 0.3954 $\mathcal{L}_{\mathrm{total}}*$ 57.96 61.35 61.81 78.54 81.52 63.26
- p-tuning (encoder) 0.1097 $\mathcal{L}_{\mathrm{total}}*$ 58.35 63.65 64.29 79.01 77.49 74.30
- prompt-tuning (embedding) 0.0138 $\mathcal{L}_{\mathrm{total}}*$ 54.07 60.30 60.56 78.34 88.18 43.00
  • ::*:∗ : To validate the effectiveness of our proposed multi-expert structure in mitigating different stereotypes, all baselines of conventional structures (e.g., LoRA) are compared using the same learning objectives. Since no expert module is involved in baselines, $\mathcal{L}_{\mathrm{expert}}$ is excluded in $\mathcal{L}_{\mathrm{total}}$.
Performance. Following the recent implementation of LLM-based recommendations [11, 29, 42], we adopt the area under the receiver operating characteristic (AUC) for performance evaluation. Meanwhile, considering the imbalanced classes (i.e., stereotype groups) in real-world recommendation datasets, the model may frequently make incorrect predictions for minority classes, yet still achieve a high AUC. Therefore, we further compare the precision and recall for performance evaluation.
VII-A3 Baselines
@@
Fairness-oriented Methods. IFairLRS [14] proposes a reweighting strategy to mitigate biases stemming from unbalanced groups, where sample weights are applied to the loss of instruction-tuning samples of LLM-RS. UP5 [11] introduces a counterfactually fair prompting method, which masks sensitive user information by prompt tuning with a discrimination loss. FaiRLLM [12] designs fairness metrics tailored to LLM-RS by mitigating the divergence of similarity metrics of recommendations against sensitive attributes.
LLM-RS Paradigms. To validate the effectiveness of our proposed multi-expert structure in mitigating different stereotypes, we compare MoS to conventional training paradigms of LLM-RS. The baselines include full-model tuning and parameter-efficient fine-tuning (PEFT) paradigms, such as prompt tuning [36], p-tuning [43], and LoRA [44]. In particular, LoRA adds extra trainable adapter weights that are merged with LLM parameters. P-tuning trains a prompt encoder to generate learnable prompt embeddings in the LLM input. Similarly, prompt-tuning directly updates the prompt embeddings. For a fair comparison between MoS (i.e., multi-expert) and baselines (i.e., non-expert), the learnable modules of each PEFT baseline are trained using the same learning objectives.
VII-A4 Implementation Details
To implement LLM-RS in a generative paradigm, the recommendation task is formulated into prompts. In particular, the personalized preference of users is indicated based on median ratings, which are 3 and 5 in MovieLens and BookCrossing datasets, respectively. An example prompt is provided below:
@@ VII-A4 Implementation Details
Our proposed methods are implemented based on HuggingFace and PyTorch. For a fair comparison across all baselines, we employ a widely-used lightweight LLM, i.e., T5 [45], with encoder-decoder structures as the backbone model of LLM-RS. As for the proposed MoS framework, we employ linear stereotype-wise routing models (i.e., router and reweighting layer) and $4$ prompt-based experts, each with a length of $5$ tokens for eliciting different stereotypes in personalized recommendations. In particular, we optimize the aforementioned proposed models with Adafator [46], where the learning rates for prompt-based baselines (i.e., p-tuning, prompt tuning, and MoS) and adapter-based baselines (i.e., TALLRec and LoRA) are set to be $0.5$ and $0.005$, respectively.
VII-B Performance Comparison
VII-B1 Comparison of Fairness-oriented Methods
As shown in Table IV, we compare the fairness and performance between our proposed MoS and existing fairness-oriented methods of LLM-RS. Overall, MoS significantly outperforms the current single-sided fairness baselines (i.e., user-side and item-side) to mitigate stereotypes in LLM-RS. In the meanwhile, MoS achieves slightly better or comparable recommendation performance compared to each baseline.
@@ VII-B1 Comparison of Fairness-oriented Methods
VII-B2 MoS vs. Conventional LLM-RS Paradigms
To demonstrate the effectiveness of MoS to learn unbiased representations utilizing multiple stereotype-relevant experts, we design baselines by replacing the MoS framework (i.e., multiple experts with stereotype-wise routing) with conventional PEFT methods to train LLM-RS under various setups of learning objectives, as shown in Table V.
Comparing the fairness performance of training LLM-RS with the proposed MoS framework to conventional PEFT paradigms (i.e., prompt tuning, p-tuning, and LoRA), it can be noticed that MoS consistently outperforms all baselines by 5%-14% in average fairness and achieves comparable AUC performance between 0.1% and 18% compared to baselines, even with higher trainable parameters of 331%-1194%. Notably, significant trade-offs between fairness and performance can be noticed in some conventional LLM-RS paradigms, while MoS maintains the best or second best scores of both fairness and performance in most situations.
TABLE VI: Results of ablation studies on MoS components. We report the average fairness ($\bm{SF}$) and performance (AUC) over three independent runs with random seeds.
@@
VII-C Ablation Study
VII-C1 MoS Components
To assess the influence of each key component, we conducted ablation experiments on the effectiveness of MoS with separately eliminated components, as shown in Table VI. In particular, we compare the stereotype-wise routing component between top-1, top-2, and stochastic routing strategies, as illustrated in Eq. (8). Notably, the top-1 setting outperforms top-2 and stochastic settings in terms of both fairness and performance. We speculate that the top-2 routing strategy potentially encodes different stereotype information to the same experts. In addition to the routing strategy, we compare the effectiveness of MoS without the reweighting layer and corresponding learning objectives $\mathcal{L}_{div}$. Despite comparable results in performance, the proposed components significantly improve the fairness by 3%-16%.
@@ VII-C1 MoS Components
Figure 5: Results of ablation studies on user-side and item-side stereotypes. The detailed statistics of users $u$, items $v$, and groups $G_{1},G_{2}$ can be found in Table III.
VII-C2 User-side vs. Item-side Stereotypes
Since a stereotype group contains both users and items in LLM-RS, we further delve into the effectiveness of MoS by separately comparing user-side stereotype and item-side stereotypes, as shown in Figure 5. By comparing between red/blue and yellow/green bars, it can be observed that MoS can achieve consistent fairness and performance between users and items of different stereotype groups. However, the fairness and performance of user stereotype group $u\in G_{1}$ (i.e., red and yellow bars) significantly exceed that of $u\in G_{2}$ (i.e., blue and green bars). One likely reason is that the amount of user-item interactions is dominated by $u\in G_{1}$ due to the unequal distribution of user groups in real-world recommendation datasets, as illustrated in Table III, leading to more sampled training data.
@@
Figure 6: Results of ablation studies on the threshold of stereotype measurement based on Z-scores.
VII-C3 Threshold for Stereotype Measurement
To validate the threshold of stereotype measurement based on Z-scores, we compare the recommendation performance of target items below and above the designed threshold. As revealed by our preliminary experiments in Table I, items below the threshold are weakly biased to any stereotype group, indicating a lower degree of stereotype (i.e., measured by $d_{v\in G}$) that potentially downgrades LLM-RS performance. It is worth noting that the comparison of fairness metrics is ignored since these items (i.e., below the threshold) are not involved in any stereotype group. As shown in Figure 6, the recommendation performance of items below the threshold notably exceeds that of items above the threshold, implying that Z-scores can identify an effective threshold for stereotype measurement.
VIII CONCLUSION
@@
This study investigates the unique characteristics of LLM stereotypes in recommendation fairness, which simultaneously involves user and item groups (i.e., two-sided groups). In this paper, we propose a new variant of fairness, extending the concept of calibrated fairness, to evaluate the influence of LLM stereotypes on a two-sided (i.e., user-to-item) recommendation fairness. In other words, the stereotypical linguistic associations, encoded by LLMs, between users and items should not amplify a user’s preference for items in recommendation tasks. To mitigate unfairness due to stereotypes in LLM-RS, a novel framework called MoS is proposed along with effective learning objectives. In particular, we develop multiple stereotype-relevant experts to capture different stereotypes in textual user and item information, where an insightful stereotype-wise routing strategy is designed to learn unbiased representations against different stereotypes. Through comprehensive experiments on recommendation datasets under various fairness settings, we demonstrated the effectiveness of the proposed methods in addressing stereotype-aware fairness of LLM-RS. As for future work, further investigation might shed light on the individual-level fairness against diverse combinations of stereotypes in LLM-based recommendations.